OpenStudy (anonymous):

can someone please show me this step by step...until you get the answer please? y=-2/3x+5 plug in -1 for x

OpenStudy (anonymous):

Type in yours step by step, we will tell you if you made a mistake.

OpenStudy (dumbcow):

y=17/3

OpenStudy (anonymous):

-2/3*-1+5 -2/3*-1=2/3 2/3+5= stumped here

OpenStudy (anonymous):

Think of 5 as equal to 15/3 2/3 + 15/3 = 17/3 Practice adding fractions.

OpenStudy (anonymous):

how are you getting 15/3 this is where I am confused

OpenStudy (anonymous):

It is a technique you pick up when you have a lot of experience in math. If the numbers are simple, convert them so that they have same denominator (the thing at the bottom); then you could just add the top and keep the bottom.

OpenStudy (anonymous):

why not just make it 7/3

OpenStudy (anonymous):

An equation, or an equality has to be true or real. 5 is not the same as 7/3.

OpenStudy (anonymous):

ok, thank you...I am going back to regular addition and subtraction..my teacher hasn't said anything about this. now I am really confused.

OpenStudy (yuki):

most people don't like math because of fractions. If you need any help, just ask us.

OpenStudy (anonymous):

I see, you added 5 to 2 to get 7/3. But you can't just add them 2/3 is like two thirds of one pie. You want to add 2/3 (or two thirds of a pie) to 5 or five pies. If you have two thirds of a pie and five pies, you really have five pies and two thirds of a pie. 17/3 is really 5 (2/3).

OpenStudy (anonymous):

so basically it is an mixed fraction 5 2/3 cahnged to an improper fraction yes? 17/3

OpenStudy (anonymous):

Right, that's it.

OpenStudy (anonymous):

wow, that is so much easier..thank you so much..now I can make a 100 on my test..thanks again.